# MSFMamba: Multi-Scale Feature Fusion State Space Model for Multi-Source Remote Sensing Image Classification

### v1

@@ IV-C Performance Comparison
We compare the proposed MSFMamba with existing state-of-the-art methods, including TBCNN [36], FusAtNet [38], $S^{2}$et [39], DFINet [55], AsyFFNet [56], ExViT [42], HCT [41] and MACN [43]. TBCNN [36] investigates the classification fusion of hyperspectral imagery and data from other sensors using a two-branch convolution neural network. FusAtNet [38] presents a multi-source classification framework by utilizing a self-attention mechanism for spectral features and a cross-attention approach for spatial features. $S^{2}$et [39] proposes a spatial-spectral enhancement module for cross-modal information interaction, which effectively facilitates the feature interaction between multi-source data. DFINet [55] uses a depth-wise cross attention module to extract complementary information from multisource feature pairs. AsyFFNet [56] is a multi-source data classification method based on asymmetric feature fusion, which employs weight-share residual blocks for feature extraction and a feature calibration module for the spatial-wise multi-source feature modeling. ExViT [42] utilizes parallel branches of position-shared Transformer extended with separable convolution modules to process multi-modal image patches. In MACN [43], mixing self-attention and convolution Transformer layer is proposed to extract local and global multi-scale feature perception.
TABLE V: Influence of MSpa-Mamba, Spe-Mamba and Fus-Mamba on classification results of MSFMamba.
MSpa-Mamba Spe-Mamba     Fus-Mamba      Berlin Augsburg Houston2018
@@
1) Results on the Berlin Dataset. The classification performance of different methods on the Berlin dataset are shown in Table II. The corresponding classification maps are shown in Fig. 7. The proposed MSFMamba achieves the best OA of 77.11% and the best Kappa of 65.57%. It is evident that the proposed MSFMamba has strong multi-scale spatial contextual modeling capability from the sequence modeling perspective. From the visualized results, it is evident that the classification map generated by our MSFMamba is closer to the ground truth and more accurately in the details of local textures.
2) Results on the Augsburg Dataset. Table III illustrates the quantitative results of different methods on the Augsburg dataset. The corresponding classification maps of different methods are shown in Fig. 8. Benefiting from extraction of spatial-spectral semantic features with the sequential modeling mechanism, the proposed MSFMamba exhibits the most clear boundaries, fewer misclassified regions. From Table III, it can be observed that our MSFMamba improves upon the second-best algorithm by 0.16% in OA, 1.91% in AA, and 0.33% in Kappa. The excellent performance improvements by our MSFMamba verifies the effectiveness of sequential scanning model for multi-source data classification.
3) Results on the Houston2018 Dataset. Table IV illustrates the quantitative results of various models on the Houston2018 dataset. Fig. 9 shows the corresponding classification maps of different methods. The proposed MSFMamba produces a well-defined classification map, in which the most noise regions are removed. From the visualized classification results in Fig. 9, the proposed MSFMamba is observed to be the closest to the ground truth, with fewer misclassified regions. The experimental results demonstrate that sequential scanning model is especially effective for HSI and LiDAR data joint classification.
@@
TABLE VI: Comparative Analysis of Model Parameters and FLOPs on the Augsburg Dataset
@@
IV-D Ablation Study
To evaluate the effectiveness of the key components of our proposed MSFMamba model, we conducted ablation studies on three datasets. These studies aim to explore the impacts of MSpa-Mamba, Spe-Mamba, and Fus-Mamba. Table V shows the effects of different modules on the classification results. The results indicate that using MSpa-Mamba as a fundamental module achieves 74.88%, 89.05%, and 89.56% for OA on three datasets, respectively. The inclusion of Spe-Mamba further enhances the classification accuracies to 76.00%, 90.60%, and 91.44%. This demonstrates that Spe-Mamba plays a crucial role in classification, highlighting the importance of spectral information for joint classification of HSI and SAR/LiDAR data. Using MSpa-Mamba and Fus-Mamba achieves classification accuracies of 75.88%, 90.55%, and 91.76%, respectively. This indicates that better fusion of multi-source data is crucial for classification tasks. Finally, the combination of all three components achieves the best performance on three datasets. It demonstrates that the MSpa-Mamba, Spe-Mamba, and Fus-Mamba all contributes to the enhancement of classification accuracy.
We conducted t-SNE visualizations of the three components on the Berlin dataset for our MSFMamba, and the results are illustrated in Fig. 10. It can be observed that with only MSpa-Mamba, the features of different classes are less compact. With all the three modules, the feature representations display the most distinct, well-defined clusters compared with the others. It is evident that sequential scanning model generate more robust feature representations for multi-source data classification.
IV-E Analysis of Computational Complexity
We investigate the computational complexity of the proposed MSFMamba, with an emphasis on model parameters and FLOPs. Table VI provides the detailed model parameters and FLOPS of different methods on the Augsburg dataset. As can be observed that the proposed MSFMamba requires fewer parameters than most of the existing methods. The FLOPS of our MSFMamba is the lowest, which indicate that our model has an advantage in computational efficiency. It is beneficial for deploying and running the proposed MSFMamba on edge devices such as drones and mobile devices.
@@ V Conclusions and Future Work
In this paper, we present a novel MSFMamba network to tackle the multi-source remote sensing data joint classification task. SSM-based blocks are designed to leverage the advantages of linear complexity and large receptive field. Specifically, we design MSpa-Mamba block for spatial feature extraction. The multi-scale strategy is incorporated to minimize the computational redundancy and alleviate feature redundancy of SSM. In addition, Spe-Mamba block is incorporated for spectral feature extraction from the hyperspectral data, which is essential for HSI and LiDAR/SAR data fusion. Moreover, we design Fus-Mamba block for multi-source feature fusion. The original Mamba is extended to accommodate dual inputs, and cross-modal feature interaction is enhanced. The proposed MSFMamba achieves state-of-the-art performance on three multi-source remote sensing datasets, highlighting the effectiveness of MSFMamba.
In the future, we will explicitly explore the hybrid framework of Mamba and Transformer to further improve the long-range feature modeling capacity. Moreover, for a more comprehensive multi-source remote sensing data interpretation, it may be necessary to study the interpretability of the Mamba model for mulit-source remote sensing data applications.

### v2

@@ IV-C Performance Comparison
We compared the proposed MSFMamba with existing state-of-the-art methods, including Two-Branch CNN(TBCNN) [49], FusAtNet,Spatial-Spectral Cross-Modal Enhancement Network($S^{2}$et) [52], Depthwise Feature Interaction Network(DFINet) [69], Asymmetric Feature Fusion Network(AsyFFNet) [70], ExViT, Hierarchical CNN and Transformer(HCT) [54] and Mixing self-Attention and Convolution Network(MACN) [56]. TBCNN [49] investigates the classification fusion of hyperspectral imagery and data from other sensors using a two-branch convolution neural network. FusAtNet [51] presents a multi-source classification framework by utilizing a self-attention mechanism for spectral features and a cross-attention approach for spatial features. $S^{2}$et [52] proposes a spatial-spectral enhancement module for cross-modal information interaction, which effectively facilitates the feature interaction between multi-source data. DFINet [69] uses a depth-wise cross attention module to extract complementary information from multisource feature pairs. AsyFFNet [70] is a multi-source data classification method based on asymmetric feature fusion, which employs weight-share residual blocks for feature extraction and a feature calibration module for the spatial-wise multi-source feature modeling. ExViT [55] utilizes parallel branches of position-shared Transformer extended with separable convolution modules to process multi-modal image patches. In MACN [56], mixing self-attention and convolution Transformer layer is proposed to extract local and global multi-scale feature perception. The classification performance of various methods on multiple datasets is presented in Tables II, III, IV, and V. These tables compare the recall rates of different methods for specific classes, providing a detailed evaluation of their performance.
TABLE VI: Influence of MSpa-Mamba, Spe-Mamba and Fus-Mamba on classification results of MSFMamba.
MSpa-Mamba Spe-Mamba     Fus-Mamba      Berlin Augsburg Houston2018 Houston2013
@@
1) Results on the Berlin Dataset. The classification performance of various methods on the Berlin dataset is presented in Table II, with corresponding classification maps shown in Fig. 8. Our proposed MSFMamba achieves the highest OA of 76.92%, an AA of 64.88%, and a Kappa coefficient of 64.88%. These results highlight the model’s robust multi-scale spatial contextual modeling capabilities, enabled by its sequential modeling approach. A closer analysis reveals that MSFMamba excels in distinguishing fine-grained classes, particularly Industrial Area, Low Plants, and Allotment, where capturing intricate textures and boundary details is critical. The model effectively handles these challenges, resulting in more accurate classification. In the visualized results, particularly in the magnified sections, it is evident that MSFMamba produces classification maps much closer to the ground truth. While other methods tend to misclassify areas like Commercial Area, MSFMamba more accurately captures these regions, showcasing its ability to preserve finer details and textures in challenging areas.
2) Results on the Augsburg Dataset. Table III presents the quantitative results of various methods on the Augsburg dataset. A closer analysis of class-wise performance reveals that MSFMamba achieves notable improvements in categories such as Low Plants, Allotment and Water, which are characterized by complex spectral-spatial patterns. Our MSFMamba outperforms the second-best algorithm by 0.22% in OA, 0.12% in AA, and 0.39% in Kappa. The corresponding classification maps, shown in Fig. 9, further demonstrate the advantage of MSFMamba. By effectively extracting spatial-spectral semantic features through its sequential modeling mechanism, MSFMamba produces clearer boundaries and significantly reduces misclassified regions compared to other methods.
3) Results on the Houston2018 Dataset. Table IV presents the classification performance of various methods on the Houston2018 dataset, with Fig. 10 showing the corresponding classification maps. MSFMamba delivers a highly refined classification map, with most noisy regions significantly reduced, resulting in cleaner and more accurate segmentation. On the Houston2018 dataset, MSFMamba demonstrates exceptional class-specific performance, especially in categories like Stressed grass, Evergreen trees, Non-residential buildings and Roads, where it surpasses other methods by better capturing spatial structures and avoiding common misclassifications. The sequential scanning approach employed by MSFMamba proves to be highly effective for HSI and LiDAR data fusion, leading to more accurate and consistent class predictions, as evidenced by the minimal noise and misclassified areas in the visualized results.
4) Results on the Houston2013 Dataset. Table V summarizes the classification results of various approaches on the Houston2013 dataset, while Fig. 11 illustrates the respective classification maps. MSFMamba produces a well-refined classification output, achieving more precise segmentation and significantly minimizing noise in most areas, resulting in a clearer result overall. On the Houston2013 dataset, MSFMamba exhibits outstanding performance on specific classes, particularly in categories like Trees, Commercial, Road and Highway. It outperforms other methods by accurately capturing spatial structures and mitigating common misclassification errors.
5) Practical Implications. The experimental results across all datasets highlight MSFMamba’s applicability in addressing real-world challenges such as land cover classification and urban infrastructure planning. For example, its high classification accuracy on datasets like Berlin and Augsburg makes it valuable for urban planning and ecological monitoring, while its robust performance on Houston2018 and Houston2013 datasets supports transportation network analysis and city development projects.
@@
IV-D Ablation Study
Table VI presents the classification performance of different module combinations on the four datasets. The results reveal a clear trend: the inclusion of each module consistently improves the OA. MSpa-Mamba serves as the fundamental module, achieving OA values of 74.88%, 89.05%, 89.56%, and 90.27% on the four datasets, respectively. Its strong spatial feature extraction capability lays a solid foundation for classification tasks. Adding Spe-Mamba further enhances performance to 76.00%, 90.60%, 91.44%, and 91.55%. This significant improvement underscores the importance of spectral feature extraction in capturing the unique patterns of HSI and SAR/LiDAR data. The combination of MSpa-Mamba and Fus-Mamba yields accuracies of 75.88%, 90.55%, 91.76%, and 91.82%. This demonstrates the critical role of effective multi-source data fusion in handling heterogeneous datasets. When all three components are combined, the model achieves the best performance across four datasets, with substantial noise reduction and more precise class boundaries. This demonstrates the complementary nature of the modules: MSpa-Mamba and Spe-Mamba provide robust spatial-spectral feature extraction, while Fus-Mamba effectively integrates these features, maximizing classification accuracy.
We conducted t-SNE (t-distributed Stochastic Neighbor Embedding) visualizations of the three components on the Berlin dataset for our MSFMamba, and the results are illustrated in Fig. 12. t-SNE is a dimensionality reduction technique widely used to visualize high-dimensional data by projecting it into a lower-dimensional space while preserving its local structure. It can be observed that with only the MSpa-Mamba, the features of different classes are less compact. With all the three modules, the feature representations display the most distinct and well-defined clusters compared to the others. It is evident that the sequential scanning model generate more robust feature representations for multi-source data classification.
TABLE VII: Experimental Results of the Multi-Scan Strategy on the Berlin Dataset.
   Path number for downsampling OA AA Kappa
0 74.71 61.15 61.99
1 75.44 62.85 62.93
2 76.92 64.88 64.88
3 75.30 62.50 62.91
4 74.31 61.79 61.99
IV-E Analysis of Multi-Scale Spatial Mamba
Table VII presents the impact of varying the number of downsampling paths on the classification performance for the Berlin dataset. The results show that the best performance is achieved when two paths retain the original resolution and two paths are downsampled, yielding OA, AA, and Kappa values of 76.92%, 64.88%, and 64.88%, respectively. This configuration balances the reduction of redundant features with the preservation of critical details, demonstrating the effectiveness of the multi-scale scanning approach. In contrast, configurations with three or more downsampling paths lead to a performance decline, with OA dropping to 75.30% and 74.31%, highlighting the trade-off between feature reduction and information loss.
This analysis emphasizes the importance of integrating original-resolution and downsampled paths to capture both fine-grained details and broader spatial patterns. The consistent improvement in classification metrics from 0 to 2 downsampling paths underscores the contribution of this multi-scale strategy in handling complex spectral-spatial patterns in the Berlin dataset. These results demonstrate that an optimal balance in downsampling is critical for achieving robust performance in remote sensing classification tasks.
Figure 13: t-SNE Visualization of Original and Fused HSI-SAR Features in Fusion Mamba.(a) Original HSI Features. (b) Original SAR Features. (c) Fused HSI Features after Fusion Mamba. (d) Fused SAR Features after Fusion Mamba.
TABLE VIII: Impact of Scanning Routes on Classification Performance
Scanning Routes OA(%) AA(%) Kappa
2 75.28 63.22 62.28
4 76.92 64.88 64.88
6 74.98 63.45 63.39
TABLE IX: Comparative Analysis of Model Parameters, FLOPS and Inference time on the Augsburg Dataset
@@ IV-E Analysis of Multi-Scale Spatial Mamba
Inference time (s) 0.2469 0.2489 0.3182 0.2924 0.3271 0.7524 0.1917 0.1747
IV-F Analysis of Fusion Mamba Effects
We aim to understand how the Fusion Mamba block enhances feature representation in multi-source remote sensing data by using t-SNE visualizations. The comparison of features before and after fusion provides insights into the strengths of the Fusion Mamba block in integrating diverse data modalities.
Fig. 13 presents the t-SNE visualizations of HSI and SAR features before and after applying the Fusion Mamba block: Fig. 13(a) shows the original HSI features before fusion. Fig. 13(b) presents the original SAR features before fusion. In both Fig. 13(a) and Fig. 13(b), the separability between different classes is relatively poor, with noticeable overlap between clusters. This suggests that the original features alone struggle to clearly discriminate between certain categories.
Fig. 13(c) demonstrates the fused HSI features after the Fusion Mamba block is applied. Fig. 13(d) displays the fused SAR features after the fusion process. The visualizations in Fig. 13(c) and Fig. 13(d) show a significant improvement in class distinction, with clearer and more compact clusters. This indicates that the Fusion Mamba block successfully integrates features from different modalities, enhancing the overall representation and improving the discriminative power of the fused features.
IV-G Analysis of Scanning Route Numbers
The number of scanning routes is an important parameter in our MSFMamba model, as it directly impacts the diversity of spatial feature representations. In this section, we evaluate the effect of different numbers of scanning routes on classification performance and computational efficiency. Specifically, we compared configurations with 2, 4, and 6 scanning routes using the Berlin dataset. In the case of 2 routes, row-wise and column-wise scanning is utilized. In the case of 4 routes, row-wise, column-wise, reverse row-wise, and reverse column wise scanning is used. To achieve 6 routes, two diagonal scanning routes (top-left to bottom right, and top-right to bottom-left) are added.
From Table VIII, we observe that the classification performance improves with the addition of reverse scanning directions. The model with 4 scanning routes achieved the highest performance, indicating that adding more scanning routes enhances the spatial feature diversity and improves the model’s ability to differentiate between classes. However, the configuration with 6 scanning routes, while providing more feature diversity, results in slightly lower performance. This could be due to the increased complexity, which may introduce more noise into the features or lead to overfitting on certain classes. Our analysis shows that using 4 scanning routes offers the best classification performance in terms of accuracy. These results suggest that adding more scanning routes does not always result in a proportional improvement in performance and may even lead to diminishing returns.
IV-H Analysis of Computational Complexity
To provide a comprehensive computational complexity analysis, we evaluated the model parameters, floating point operations per second(FLOPS), and inference time of various methods on the Augsburg dataset, as shown in Table IX. Specifically, we measure the total number of learnable parameters in millions (M), the computational cost in gigaflops (G), and the inference time in seconds (s). Our proposed MSFMamba model exhibits the lowest FLOPs and inference time, confirming its computational efficiency compared to other methods. This advantage highlights MSFMamba’s suitability for real-time applications, especially on resource-constrained edge devices like drones and mobile platforms. Furthermore, MSFMamba requires significantly fewer parameters than most state-of-the-art methods, contributing to reduced memory requirements and facilitating efficient deployment.
@@ V Conclusions and Future Work
This paper introduced MSFMamba, a novel network designed for the joint classification of multi-source remote sensing data. The network utilized SSM-based blocks to achieve linear complexity and a large receptive field. The MSpa-Mamba block efficiently extracted spatial features using a multi-scale strategy, reducing computational and feature redundancies. The Spe-Mamba block enhanced spectral feature extraction, enabling effective HSI and LiDAR/SAR fusion, while the Fus-Mamba block facilitated robust cross-modal feature integration by extending the original Mamba design to support dual inputs. Experiments on four datasets demonstrated that MSFMamba achieved state-of-the-art performance, highlighting its robustness, efficiency, and potential for real-world applications.
However, the model has certain limitations, such as reduced accuracy for imbalanced classes and limited interpretability in critical applications like urban planning and disaster management. Addressing these challenges requires exploring techniques such as dynamic sampling, advanced loss functions, and methods to enhance explainability. Future work will focus on developing hybrid models that combine Mamba with Transformer architectures to further improve long-range feature modeling and multi-modal data fusion. Additionally, methods for improving the interpretability of MSFMamba will be explored, particularly in terms of how it processes and integrates the contributions of different data modalities.Investigating feature importance and feature selection techniques will also be part of the future research, ensuring that the model is both efficient and transparent in its operations.